OpenStudy (chris215):

To three decimal places, find the value of the first positive x-intercept for the function f(x) = 2cos(x + 3)

OpenStudy (anonymous):

\[\frac{5\pi}{2}-2\] is what i was trying to suggest in the last post

OpenStudy (anonymous):

okk actually \[\frac{5\pi}{2}-3\]

OpenStudy (chris215):

1.712

OpenStudy (chris215):

but then i get -1.429 and thats the wrong answer

OpenStudy (merchandize):

\[x=\pi \div2-3\]

OpenStudy (chris215):

i got the same ansewer tho

OpenStudy (merchandize):

do you have answer for that?

OpenStudy (chris215):

-1.429..

OpenStudy (chris215):

im confused that answer isn't correct though bc i checked w my teacher and the answer is supposed to be postive thats all i knwo

OpenStudy (merchandize):

I need to check again..there can be a mistake